OpenStudy (anonymous):

Solve for the roots in the following equation. (x^4 + 5x^2 - 36)(2x^2 + 9x - 5) = 0

geerky42 (geerky42):

(x^4 + 5x^2 - 36)(2x^2 + 9x - 5) = 0 means that either x^4 + 5x^2 - 36 = 0 or 2x^2 + 9x - 5 = 0

geerky42 (geerky42):

for x^4 + 5x^2 - 36, you can set x^2 = u so it will be u^2+5u-36 = 0, then you can use quadratic formula to solve for u.

geerky42 (geerky42):

For 2x^2 + 9x - 5 = 0, use quadratic formula too

OpenStudy (triciaal):

you can never go wrong with the formula 2x^2 + 9x -5 = 0 can be factored as (x + 5)(2x - 1)= 0 @Power2Knowledge are you sure in the first expression its 36 and not 6 so it could be easily factored (u + 6)(u - 1)=0